OpenStudy (anonymous):

Solve the system of linear equations below. 2x + 3y = 2 x + 6y = 4

OpenStudy (anonymous):

by using simultineous equation multiply the first equation by (-2) -4x-6y=-4 x+6y=4, then add up them, -3x=0; x=0 then the value of y is 2(0)+3y=2 y=2/3

OpenStudy (theviper):

first find the value of x in any eq than put the value in other eq than find x and y:)